OpenStudy (anonymous):

A square support unit in a television is made with a side measuring 5 centimeters. A new model being designed for the upcoming year will have a larger square with a side measuring 5.2 centimeters. By how much will the area of the square be increased?

OpenStudy (anonymous):

\[A1 = (5)^2 = 25\]\[A2 = (5.2)^2 = 27.04\] so the increase in the area is : A(increased) = 27.04 - 25 = 2.04 so it increased by 2.04 ^_^

OpenStudy (anonymous):

Sweet.. yes that was one of the answers...

OpenStudy (anonymous):

yay :)